# Boundary transfer matrices arising from quantum symmetric pairs

### v1

@@ Abstract.
The first author is partially supported by an INdAM Project Grant 2024. Both authors are partially supported by the International Scientists Program of the Beijing Natural Sciences Foundation (grant number IS24003).
@@ 1.1. Quantum symmetric pairs
Motivated by questions in harmonic analysis, quantum deformations of the homogeneous spaces $G/G^{\Theta}$ were estensively studied in the 1990s. Here $G$ is a connected complex Lie group, while $G^{\Theta}$ is the subgroup of fixed points with respect to an involutive automorphism $\Theta\colon G\to G$. Let $\theta=d\Theta$ be the involutive automorphism of the corresponding Lie algebra ${\mathfrak{g}}$ and consider the fixed-point subalgebra ${\mathfrak{g}}^{\theta}$.
@@
In parallel, G. Letzter’s introduced one-sided coideal deformations of $U({\mathfrak{g}}^{\theta})$ for any symmetric pair $({\mathfrak{g}},{\mathfrak{g}}^{\theta})$, see [Let97, Let99, Let02]. Their definition is given directly in terms of the generators of $U_{q}({\mathfrak{g}})$. This construction was extended by S. Kolb at the level of symmetrizable Kac-Moody algebras in [Kol14], recovering as a special case the twisted quantum loop algebras, also known as twisted q-Yangians, see [MRS03, CGM14]. Quantum deformations $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ of a more general class of subalgebras ${\mathfrak{k}}\subseteq{\mathfrak{g}}$ which are not necessarily fixed-point subalgebras, which we refer to as pseudo-fixed-point subalgebras, were obtained in [RV20, RV22]. We refer to all such pairs $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ as quantum (pseudo-)symmetric pairs (QSP), and to $U_{q}(\mathfrak{k})$ as a QSP subalgebra.
@@ 1.2. Universal K-matrices
In the works [NS95, Nou96, MRS03, KS09, CGM14] the QSP subalgebras were defined in terms of a K-matrix, i.e.,  a matrix solution to a reflection equation, using an approach akin to the R-matrix presentation for quantum groups outlined in [FRT90]. Conversely, in the Letzter-Kolb formalism the QSP subalgebra $U_{q}({\mathfrak{k}})\subseteq U_{q}({\mathfrak{g}})$ naturally gives rise to a universal solution of a (generalized) reflection equation. In [BW18a], H. Bao and W. Wang introduced a twisted $U_{q}(\mathfrak{k})$-intertwiner, called quasi K-matrix, and proved its existence for a family of QSPs extending $U_{q}(\mathfrak{sl}_{N})$. Their construction has been generalized across the range of QSPs, see [BK19, RV20, AV22a, WZh23].
@@
In [AV22a], we introduced the notion of a cylindrical structure on $(A,B)$ where $A$ is a quasitriangular bialgebra and $B\subseteq A$ is a right coideal subalgebra. It consists of a suitable algebra automorphism $\psi:A\to A$ and a twisted $B$-intertwiner $K\in A^{\times}$, called a basic K-matrix, satisfying a twisted reflection equation with respect to the universal R-matrix $R$ of $A$. Cylindrical structures can be gauge transformed by elements in $A^{\times}$. This gauge action can be used to reduce either $K$ or $\psi$ to a more convenient form.
@@
This approach has been instrumental in [AV22b], leading to a uniform and systematic construction of spectral K-matrices (i.e., parameter-dependent matrix solutions of the reflection equation) on finite-dimensional representations of quantum affine algebras. The key idea is that the twist automorphism $\psi$ can be used to accomplish the inversion of the spectral parameter, see [AV22b]. Specifically, when ${\mathfrak{g}}$ is an (untwisted) affine Lie algebra, the universal K-matrix gives rise to a formal Laurent series on any finite-dimensional module over the quantum affine algebra $U_{q}({\mathfrak{g}}^{\prime})\subset U_{q}({\mathfrak{g}})$. It satisfies a generalized reflection equation of the type considered in [FM91, KS92, Che92], which includes the original ”untwisted” reflection equation from [Che84, Skl88]. Moreover, on irreducible modules, it yields a rational solutions of the same generalized reflection equation, called a trigonometric K-matrix.
@@
By the coideal property of $B$, ${\operatorname{Mod}}(B)$ is naturally equipped with a module structure over the monoidal category ${\operatorname{Mod}}(A)$. Since ${\operatorname{Mod}}(A)$ is braided, it is natural to require that the action on ${\operatorname{Mod}}(B)$ is compatible with the braiding induced by the R-matrix. This is equivalent to the existence of an element ${{\mathbb{K}}}\in B\otimes A$ which satisfies similar relations as $K$, and it is therefore referred to as a tensor K-matrix. The universal framework encoding the action of ${{\mathbb{K}}}$ has been developed in [DKM03, Enr07, Bro13, BZBJ18], with further developments and applications to the theory of quantum symmetric pairs in [DCNTY19, Kol20, KY20, LBG23, Lem23, AV24]. Roughly, the tensor K-matrix yields a braiding on ${\operatorname{Mod}}(B)$, and induce an action of the cylindrical braid group.
@@
${{\mathbb{K}}}\coloneqq(R^{\psi})_{21}\cdot(1\otimes K)\cdot R$ (1.1)
is almost automatically a tensor K-matrix, except for the support condition ${{\mathbb{K}}}\in B\otimes A$. In the case of an arbitrary QSP, we proved in [AV24] that, for every cylindrical structure constructed in [AV22a], the operator ${{\mathbb{K}}}$ is indeed a tensor K-matrix. [^1] Motivated by this result, in this paper we include the support condition in the definition of a cylindrical structure, in contrast with [AV22a, AV22b, AV24].
@@ 1.3. Present work
In this paper, we build on the results from [AV22a, AV24] to construct a universal version of a boundary transfer matrix, see [LBG23, Lem23] and references therein, in the form of a ring homomorphism
$\mathscr{T}:[{\operatorname{Mod}}_{\sf fd}(A)]\to B$ (1.2)
depending upon the choice of tensor K-matrix ${{\mathbb{K}}}\in B\otimes A$. Here, $[{\operatorname{Mod}}_{\sf fd}(A)]$ denotes the Grothendieck ring of finite-dimensional $A$-modules.
Analogous to the case of the transfer matrices arising from a universal R-matrix, it would be natural to define $\mathscr{T}$ by allowing the second leg of ${{\mathbb{K}}}\in B\otimes A$ to act on a finite-dimensional $A$-module $V$ and then taking the trace over $V$. This construction, however, does not quite work as ${{\mathbb{K}}}$ does not behave well with respect to tensor products in the second leg. This can be resolved by gauge-transforming the cylindrical structure with respect to a class of distinguished elements, which we term boundary gauge transformation. This modification is inspired by the theory of (non-universal) 2-boundary transfer matrices due to E. Sklyanin, see [Skl88] and also [MN91, FSHY97, Vla15].
@@
    Let $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ be a QSP of finite type equipped with the cylindrical structure from [BK19, Kol20]. When $K^{\prime}$ is the balance and $\widetilde{K}$ is the Drinfeld element, the map $\mathscr{T}$ specializes to the ring homomorphism $\bm{\Phi}$ constructed by Kolb in [Kol20, Sec. 4], and it is therefore surjective over $Z(U_{q}(\mathfrak{k}))$ (see Proposition 5.9.1).
@@
    $\mathscr{T}(z):[{\operatorname{Mod}}_{\sf fd}(U_{q}({\mathfrak{g}}))]\to{U_{q}(\mathfrak{k})}(\negthinspace(z)\negthinspace)\,.$ (1.3)
@@ 1.4. Notation
We will introduce some standard notations which will be used throughout this paper for any unital associative ${\mathbb{F}}$-algebra $A$ for any field ${\mathbb{F}}$. We denote conjugation by $a\in A^{\times}$ by $\operatorname{Ad}(a)$.
@@
:${}^{\sf t_{2}}:A\otimes\operatorname{End}_{{\mathbb{F}}}(V)\to A\otimes\operatorname{End}_{{\mathbb{F}}}(V^{\vee})$ (1.4)
@@ 1.5. Acknowledgments
Some of the new results contained in this manuscript were announced at the conference Recent Advances in Quantum Integrable Systems 2024 at Laboratoire d’Annecy de Physique Théorique. B.V. thanks the organizers for the kind invitation and hospitality.
@@ 2. Boundary transfer matrices for cylindrical bialgebras
A cylindrical structure is an extra structure placed on a quasitriangular bialgebra $A$ and a right coideal subalgebra $B$, specifically designed to encode a universal solution to a generalized reflection equation [AV22a, AV24]. In this section, we determine a sufficient condition for a cylindrical structures on $(A,B)$ to yield a ring homomorphism from the Grothendieck ring of ${\operatorname{Mod}}_{\sf fd}(A)$ to a commutative subring of $B$. This is a boundary analogue of the transfer matrix map considered in [FR99], providing a universal and generalized version of two-row transfer matrices, originally due to E. Sklyanin [Skl88].
2.1. Quasitriangular Hopf algebras and twist pairs
@@
(and hence $\epsilon\circ\psi^{-1}=\epsilon$). Equivalently, the Drinfeld twist $J$ measures the failure of $\psi$ to be an isomorphism of quasitriangular bialgebras [^3] Indeed, in the terminology introduced in [Dav07], a twist pair is the same as a twisted morphism of quasitriangular bialgebras $A^{\sf cop}\to A$. $A^{\sf cop}\to A$.
@@ 2.2. Cylindrical structures [AV22a]
This notion555 The support condition (2.8) is not included in the definition given in [AV22a]. Instead, it was considered in [Kol20], in the case $\psi$ is an automorphism of quasitriangular bialgebras, and, more generally, in [AV24]. of a cylindrical structure, given in [AV22a], naturally generalizes the formalism introduced by T. tom Dieck and R. Häring-Oldenburg in the case $\psi={\operatorname{id}}$ [tDHO98] and by M. Balagović and S. Kolb in the case $\psi\colon A\to A$ is an automorphism of quasitriangular bialgebras and $J=R_{21}^{-1}$ [BK19]. $\triangledown$
@@ 2.3. Tensor K-matrices
see [AV24, App. B] and [LBG23, Lem23] for the more general setting of right comodule algebras. ${{\mathbb{K}}}$ is referred to as a tensor K-matrix and satisfies
@@
    As an operator on modules, the action of ${{\mathbb{K}}}$ gives a braiding on the module category ${\operatorname{Mod}}(B)$ over ${\operatorname{Mod}}(A)$, and yields an action of the cylindrical braid group, see [Kol20]. However, in the setting of quantum symmetric pairs, ${\operatorname{Mod}}(A)$ is replaced by a proper subcategory ${\mathcal{C}}\subset{\operatorname{Mod}}(A)$, which may not be closed under twisting by $\psi$. In these cases, the framework of braided module categories does not apply. To remedy this, we introduced in [AV24, App. C] the notion of a boundary bimodule category.
@@
    The case $\psi={\operatorname{id}}$ was already considered in [Enr07, Bro13, Kol20]. In [AV24, App. B.5], we also highlight the similarities with the framework of weakly universal K-matrices from [KY20]. $\triangledown$
@@ 2.4. Gauge transformations and invariants
There is a natural action of the group $A^{\times}$ on the set of cylindrical structures on $(A,B)$, which we refer to as gauge transformations, see [AV22b, Sec. 3.5]. The action of $g\in A^{\times}$ is given by
${(\psi,J,K)}$${(\overline{\psi},\overline{J},\overline{K})=(\operatorname{Ad}(g)\circ\psi,(g\otimes g)\cdot J\cdot\Delta(g)^{-1},g\cdot K)\,,}$$\scriptstyle{g}$ (2.15)
and it extends to the tensor K-matrix as $\overline{\mathbb{K}}=(1\otimes g)\cdot{{\mathbb{K}}}$, see [AV24, Sec. 5.10].
@@
    The cylindrical structure $(\overline{\psi},\overline{J},\overline{K})$ is canonical in the sense that it is a gauge invariant of $(\psi,J,K)$, i.e., for any $g\in A^{\times}$, one has
[…]
@@
    In [AV24], we considered the special case $D=1$, which led us to determine a sufficient condition for (2.8). Namely, let $\xi\coloneqq\operatorname{Ad}(K^{-1})\circ\psi$ and set
@@
    Indeed, one checks easily that $(R^{\xi})_{21}\cdot R$ lies in $B_{\xi}\otimes A$, and therefore so does $(R^{\psi})_{21}\cdot K_{2}\cdot R$.

### v2

@@ Abstract.
The first author is partially supported by an INdAM Project Grant 2024 and the PRIN grant 2022HMTBLL. Both authors are partially supported by the International Scientists Program of the Beijing Natural Sciences Foundation (grant number IS24003).
@@ 1.1. Quantum symmetric pairs
Motivated by questions in harmonic analysis, quantum deformations of the homogeneous spaces $G/G^{\Theta}$ were extensively studied in the 1990s. Here $G$ is a connected complex Lie group, while $G^{\Theta}$ is the subgroup of fixed points with respect to an involutive automorphism $\Theta\colon G\to G$. Let $\theta={\sf d}\Theta$ be the involutive automorphism of the corresponding Lie algebra $\mathfrak{g}$} and consider the fixed-point subalgebra ${\mathfrak{g}}^{\theta}$.
@@
In parallel, G. Letzter introduced one-sided coideal deformations of $U({\mathfrak{g}}^{\theta})$ for any symmetric pair $({\mathfrak{g}},{\mathfrak{g}}^{\theta})$, see [Let97, Let99, Let02]. Their definition is given directly in terms of the generators of $U_{q}({\mathfrak{g}})$. This construction was extended by S. Kolb to the case where $\mathfrak{g}$} is any symmetrizable Kac-Moody algebra in [Kol14], recovering as a special case the twisted quantum loop algebras, also known as twisted q-Yangians, see [MRS03, CGM14]. Quantum deformations $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ of a more general class of subalgebras ${\mathfrak{k}}\subseteq{\mathfrak{g}}$ which are not necessarily fixed-point subalgebras, which we refer to as pseudo-fixed-point subalgebras, were obtained in [RV20, RV22]. We refer to all such pairs $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ as quantum (pseudo-)symmetric pairs (QSP), and to $U_{q}(\mathfrak{k})$ as a QSP subalgebra.
@@ 1.2. Universal K-matrices
In the works [NS95, Nou96, MRS03, KS09, CGM14] the QSP subalgebras were defined in terms of a K-matrix, i.e.,  a matrix solution to a reflection equation, using an approach akin to the R-matrix presentation for quantum groups outlined in [FRT90]. Conversely, in the Letzter-Kolb formalism the QSP subalgebra $U_{q}({\mathfrak{k}})\subseteq U_{q}({\mathfrak{g}})$ naturally gives rise to a universal solution of a (generalized) reflection equation. In [BW18a], H. Bao and W. Wang introduced a twisted $U_{q}(\mathfrak{k})$-intertwiner, called quasi K-matrix, and proved its existence for a family of QSPs extending $U_{q}(\mathfrak{sl}_{N})$. This construction has been generalized across the range of QSPs, see [BW18b, BK19, RV20, AV22, WZh23].
@@
In [AV22], we introduced the notion of a cylindrical structure on $(A,B)$ where $A$ is a quasitriangular bialgebra and $B\subseteq A$ is a right coideal subalgebra. It consists of a suitable algebra automorphism $\psi:A\to A$ and a twisted $B$-intertwiner $K\in A^{\times}$, called a basic K-matrix, satisfying a twisted reflection equation with respect to the universal R-matrix $R$ of $A$. Cylindrical structures can be gauge-transformed by elements in $A^{\times}$. This gauge action can be used to bring either $K$ or $\psi$ to a more convenient form.
@@
This approach has been instrumental in [AV25a], leading to a uniform and systematic construction of spectral K-matrices (i.e., parameter-dependent matrix solutions of the reflection equation) on finite-dimensional representations of quantum affine algebras. The key idea is that the twist automorphism $\psi$ can be used to accomplish the inversion of the spectral parameter. Specifically, when $\mathfrak{g}$} is an (untwisted) affine Lie algebra, the universal K-matrix gives rise to a formal Laurent series on any finite-dimensional module over the quantum affine algebra $U_{q}({\mathfrak{g}}^{\prime})\subset U_{q}({\mathfrak{g}})$. It satisfies a generalized reflection equation of the type considered in [FM91, KS92, Che92], which includes the original ”untwisted” reflection equation from [Che84, Skl88]. Moreover, on irreducible modules, it yields rational solutions of the same generalized reflection equation, called a trigonometric K-matrix.
@@
By the coideal property of $B$, ${\operatorname{Mod}}(B)$ is naturally equipped with a module structure over the monoidal category ${\operatorname{Mod}}(A)$. Since ${\operatorname{Mod}}(A)$ is braided, it is natural to require that the action on ${\operatorname{Mod}}(B)$ is compatible with the braiding induced by the R-matrix. This is equivalent to the existence of an element ${{\mathbb{K}}}\in B\otimes A$ which satisfies similar relations as $K$, and it is therefore referred to as a tensor K-matrix. The universal framework encoding the action of $\mathbb{K}$}} has been developed in [DKM03, Enr07, Bro13, BZBJ18], with further developments and applications in [DCNTY19, Kol20, KY20, LBG23, Lem23, Xu23, AV25b]. Roughly, the tensor K-matrix yields a braiding on ${\operatorname{Mod}}(B)$, and induces an action of the cylindrical braid group.
@@
${{\mathbb{K}}}\coloneqq(R^{\psi})_{21}\cdot(1\otimes K)\cdot R,\qquad R^{\psi}\coloneqq(\psi\otimes{\operatorname{id}})(R),$ (1.1)
is almost automatically a tensor K-matrix, except for the support condition ${{\mathbb{K}}}\in B\otimes A$. In the case of an arbitrary QSP, we proved in [AV25b] that, for every cylindrical structure constructed in [AV22], the operator $\mathbb{K}$}} is indeed a tensor K-matrix. [^1] Motivated by this result, in this paper we include the support condition in the definition of a cylindrical structure, in contrast with [AV22, AV25a, AV25b].
@@ 1.3. Present work
In this paper, we build on the results from [AV22, AV25b] to provide a uniform construction of a universal boundary transfer matrix and references therein, in the form of a ring homomorphism
$\mathscr{T}:[{\operatorname{Mod}}_{\sf fd}(A)]\to B,\qquad[V]\mapsto\mathscr{T}_{V}=\mathop{\sf Tr}_{2}({{\mathbb{K}}})$ (1.2)
where ${\operatorname{Mod}}_{\sf fd}(A)]$ denotes the Grothendieck ring of finite-dimensional $A$-modules. Analogous to the case of transfer matrices arising from a universal R-matrix, we define $\mathscr{T}$ by letting the second leg of ${{\mathbb{K}}}\in B\otimes A$, to act on a finite-dimensional $A$-module $V$ and then taking the trace over $V$. Note also that in [Lem23] universal transfer matrices of this type are considered for the q-Onsager algebra.
This construction, however, does not quite work if we consider an arbitrary cylindrical structure: it does not behave well with respect to tensor products in the second leg. For a large class of bialgebras $A$, this can be resolved by gauge-transforming, resulting in a dinstinguished cylindrical structure. We call the relevant gauge transformation, which appears as a factor in the formula for the transformed $\mathscr{T}$, a boundary gauge transformation. This modification is inspired by the theory of (non-universal) 2-boundary transfer matrices due to E. Sklyanin, see [Skl88] and also [MN91, FSHY97, Vla15].
@@
    Let $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ be a QSP of finite type equipped with the cylindrical structure from [BK19, Kol20]. When $K^{\prime}$ is the balance, so $\widetilde{K}$ is the Drinfeld element, the map $\mathscr{T}$ specializes to the ring homomorphism $\bm{\Phi}$ constructed by Kolb in [Kol20, Sec. 4], and it is therefore surjective over $Z(U_{q}(\mathfrak{k}))$ (see Proposition 5.9.1).
@@
    $\mathscr{T}(z):[{\operatorname{Mod}}_{\sf fd}(U_{q}({\mathfrak{g}}))]\to{U_{q}(\mathfrak{k})}(\negthinspace(z)\negthinspace)$ (1.3)
@@ 1.4. Notation
We will introduce some standard notations which will be used throughout this paper for any unital associative $\mathbb{F}$}-algebra AA for any field $\mathbb{F}$}. We denote conjugation by $a\in A^{\times}$ by $\operatorname{Ad}(a)$.
@@
$^{\sf t_{2}}:A\otimes\operatorname{End}_{{\mathbb{F}}}(V)\to A\otimes\operatorname{End}_{{\mathbb{F}}}(V^{\vee})$ (1.4)
@@ 1.5. Acknowledgments
The authors would like to thank Pascal Baseilhac, Azat Gainutdinov, David Hernandez and Stefan Kolb for stimulating discussions. Some of the new results contained in this manuscript were announced at the conference Recent Advances in Quantum Integrable Systems 2024 at Laboratoire d’Annecy de Physique Théorique. B.V. thanks the organizers for the kind invitation and hospitality.
@@ 2. Boundary transfer matrices for cylindrical bialgebras
A cylindrical structure is an extra structure placed on a quasitriangular bialgebra $A$ and a right coideal subalgebra $B$, specifically designed to encode a universal solution to a generalized reflection equation [AV22, AV25b]. In this section, we determine a sufficient condition for a cylindrical structures on $(A,B)$ to yield a ring homomorphism from the Grothendieck ring of ${\operatorname{Mod}}_{\sf fd}(A)$ to a commutative subring of $B$. This is a boundary analogue of the transfer matrix map considered in [FR99], providing a universal and generalized version of two-row transfer matrices, originally due to E. Sklyanin [Skl88].
2.1. Quasitriangular bialgebras and twist pairs
@@
(and hence $\epsilon\circ\psi^{-1}=\epsilon$). Equivalently, the Drinfeld twist $J$ measures the failure of $\psi$ to be an isomorphism of quasitriangular bialgebras [^3] Indeed, in the terminology introduced in [Dav10], a twist pair is the same as a twisted morphism of quasitriangular bialgebras $A^{\sf cop}\to A$. $A^{\sf cop}\to A$.
@@ 2.2. Cylindrical structures [AV22]
This notion [^5] The support condition (2.8) is not included in the definition given in [AV22]. Instead, it was considered in [Kol20], in the case $\psi$ is an automorphism of quasitriangular bialgebras, and, more generally, in [AV25b]. of a cylindrical structure, given in [AV22], naturally generalizes the formalism introduced by T. tom Dieck and R. Häring-Oldenburg in the case $\psi={\operatorname{id}}$ [tDHO98] and by M. Balagović and S. Kolb in the case $\psi\colon A\to A$ is an automorphism of quasitriangular bialgebras and $J=R_{21}^{-1}$ [BK19]. $\triangledown$
@@ 2.3. Tensor K-matrices
see [AV25b, App. B] and [LBG23, Lem23] for the more general setting of right comodule algebras. $\mathbb{K}$}} is referred to as a tensor K-matrix and satisfies
@@
    As an operator on modules, the action of $\mathbb{K}$}} gives a braiding on the module category ${\operatorname{Mod}}(B)$ over ${\operatorname{Mod}}(A)$, and yields an action of the cylindrical braid group, see [Kol20]. However, in the setting of quantum symmetric pairs, ${\operatorname{Mod}}(A)$ is replaced by a proper subcategory ${\mathcal{C}}\subset{\operatorname{Mod}}(A)$, which may not be closed under twisting by $\psi$. In these cases, the framework of braided module categories does not apply. To remedy this, we introduced in [AV25b, App. C] the notion of a boundary bimodule category.
@@
    The case $\psi={\operatorname{id}}$ was already considered in [Enr07, Bro13, Kol20]. In [AV25b, App. B.5], we also highlight the similarities with the framework of weakly universal K-matrices from [KY20]. $\triangledown$
@@ 2.4. Gauge transformations and invariants
There is a natural action of the group $A^{\times}$ on the set of cylindrical structures on $(A,B)$, which we refer to as gauge transformation, see [AV22, Rmk. 8.11] and [AV25a, Sec. 3.5]. The action of $g\in A^{\times}$ is given by
$(\psi,J,K)$$(\overline{\psi},\overline{J},\overline{K})=(\operatorname{Ad}(g)\circ\psi,(g\otimes g)\cdot J\cdot\Delta(g)^{-1},g\cdot K)\,$$\scriptstyle{g}$ (2.15)
and it extends to the tensor K-matrix as $\overline{\mathbb{K}}=(1\otimes g)\cdot{{\mathbb{K}}}$, see [AV25b, Sec. 5.10].
@@
    The cylindrical structure $(\overline{\psi},\overline{J},\overline{K})$ is canonical in the sense that it is a gauge invariant of $(\psi,J,K)$: it only depends on the gauge-equivalence class of $(\psi,J,K)$. In other words, for any $g\in A^{\times}$, one has
    $(\psi,J,K)$$(\operatorname{Ad}(g)\circ\psi,(g\otimes g)\cdot J\cdot\Delta(g)^{-1},g\cdot K)$$(\overline{\psi},\overline{J},\overline{K})$$(\overline{\psi},\overline{J},\overline{K}).$$\scriptstyle{g}$$\scriptstyle{D\cdot K^{-1}}$$\scriptstyle{D\cdot(g\cdot K)^{-1}}$ (2.17)
@@
    In [AV25b], we considered the special case $D=1$, which led us to determine a sufficient condition for (2.8). Namely, let $\xi\coloneqq\operatorname{Ad}(K^{-1})\circ\psi$ and set
@@
    Indeed, (2.10) implies that $(R^{\xi})_{21}\cdot R$ lies in $B_{\xi}\otimes A$, and therefore so does $(R^{\psi})_{21}\cdot K_{2}\cdot R$.